OpenStudy (anonymous):

Solve the following problem using substitution or elimination: The sum of two numbers is 170 and their difference is 26. Be sure to show all work.

OpenStudy (anonymous):

ok 65

OpenStudy (anonymous):

x + y =170 y = (x + 26) x + (x + 26) = 170 x + x + 26 = 170 2x + 26 = 170 2x = 170 - 26 2x = 144 x = 144/2 x = 72 x + y =170 72 + y =170 y = 170 - 72 y = 98 So the first number is 72 , and the second 98